To obtain a pleasing rectangular shape, the ancient Greeks constructed buildings with a length that was about 1/6 longer than the width. If the width of Parthenon is 30.89 meters, then what is its perimeter? What is the area?
Answer by Theo(13342)

let L = length and let W = width.

a pleasing geometric shape, according to the ancient greeks, requires L = W + 1/6 * W which is equal to 6/6 * W + 1/6 * W which is equal to 7/6 * W.

you have L = 7/6 * W

if W = 30.89 meters, then L = 7/6 * 30.89 = 36.038333333 meters.

the perimeter is equal to 2L + 2W which is equal to 2 * (L + W) which is equal to 133.856666667 meters.

the area is equal to L * W which is equal to 1113.224117 square meters.